OpenStudy (kiwibyrd):

Please check my work! "Find the length of side a if angle A = 62 degrees, b = 15, and c = 21." My answer was 19.24, is that right?

OpenStudy (anonymous):

Is this a right triangle Because if it isnt I dont think we have enough info to solve it

OpenStudy (kiwibyrd):

whoops yeah its a right triangle

OpenStudy (help2001):

Equation H: y = −x + 2 Equation J: y = 3x − 4 Which of the following steps can be used to find the solution to the set of equations?

OpenStudy (anonymous):

\[\cos 62=\frac{ 15^2+21^2-a^2 }{ 2*15*21 }=\frac{ 225+441-a^2 }{ 630 }=\frac{ 666-a^2 }{ 630 }\]

OpenStudy (help2001):

MEDAL AND FAN!!!!!!!111 A set of equations is given below: Equation H: y = −x + 2 Equation J: y = 3x − 4 Which of the following steps can be used to find the solution to the set of equations? −x = 3x − 4 −x +2 = 3x −x + 2 = 3x − 4 −x + 1 = 3x + 2 @Brill WHICH OPTION IS IT :( is it D?

RhondaSommer (rhondasommer):



RhondaSommer (rhondasommer):



RhondaSommer (rhondasommer):

its about 18.5

OpenStudy (kiwibyrd):

oh, i was off by a little

RhondaSommer (rhondasommer):

that is what i would do any way :)

OpenStudy (anonymous):

\[0.47=\frac{ 666-a^2 }{ 630 },0.47*630=666-a^2\] a^2=666-296=370 a=19.24 approx.

OpenStudy (kiwibyrd):

then i was right?

OpenStudy (kiwibyrd):

oh gosh i dont know who to believe

OpenStudy (anonymous):

correct.

OpenStudy (kiwibyrd):

@surjithayer could you check a few more? I'll post them seperately, of course.

OpenStudy (anonymous):

can you show whole work?

OpenStudy (kiwibyrd):

i dont have the work for them anymore, only the answers

OpenStudy (anonymous):

because if I solve it takes too much time. ok,if possible i will check,you post.

OpenStudy (kiwibyrd):

thank you !